# Robust Multimodal Learning via Cross-Modal Proxy Tokens

### v2

@@ 5 Limitations and Future Directions
Below we list some limitations and potential future directions. First, while our approach shows improved performance with missing modalities over existing methods, missing modalities still cause performance drop. Second, we observe that certain classes are heavily dependent on a specific modality, which leads to significant performance drop in those classes when that specific modality is missing. A natural question arises as how can we design a better training strategy that encourages more uniform reliance on different modalities while maintaining overall performance in both complete and missing modality scenarios? Future work can also explore developing a general framework for explaining modality contribution and expected performance drop under different missing modality scenarios.
@@ 6 Conclusion
We presented a simple yet effective framework for utilizing pre-trained unimodal encoders for robust multimodal learning. Our proposed framework enables training of multimodal systems that offer robust performance under different scenarios of missing modalities during training and inference. To compensate for missing modalities, we propose to learn cross-modal proxy tokens (CMPTs) to approximate the missing modality class tokens from the available modality using an alignment loss during training. Instead of training the unimodal encoders from scratch, we only learn rank-1 adapters along with task-specific classifier heads. Our extensive evaluations across diverse tasks and datasets demonstrate the robustness and adaptability of our approach in both complete and missing modality scenarios. Our analysis also reveals that CMPTs can generalize well to different missing scenarios and enhance performance across most of the classes to provide and overall performance boost. Overall, our approach provides a robust and efficient solution for multimodal learning that maintains high performance in different real-world scenarios.
7 Acknowledgment
This work is supported in part by an Amazon Gift award and NSF awards 2046293 and 2406199.
Supplementary Material
S1 Related Work on Multimodal Learning
@@ 7 Acknowledgment
In this section, we provide a general overview of multimodal learning, which involves integrating information from different input sources to improve performance on downstream tasks. We can broadly categorize multimodal learning approaches into two main categories: designing specialized models and adapting existing models.
@@
Designing specialized models to effectively fuse multiple input sources is a common strategy in multimodal learning. For vision-language tasks, models such as ViLT [21], ViLBERT [27], and Visual-BERT [23] process image patches and text tokens directly within a transformer architecture for tasks like image captioning and visual question answering. MMBT [19] integrates both text and images within a BERT-based framework [7] to model cross-modal dependencies. MBT [33] leverages fusion bottlenecks for efficient modality-specific interaction, particularly in audio-video tasks. Specialized multimodal models have also been proposed for other applications, including action recognition [51, 56], segmentation [53, 49], and sentiment analysis [51, 43]. While these models excel in specific tasks, extending them to new modalities or domains often proves challenging and resource-intensive, as it requires retraining or significant modification of the model architecture.
@@
Adapting existing models offers an alternative approach to multimodal learning, focusing on the use of pre-trained unimodal models and adapting them for multimodal tasks. One such strategy is gradient modulation, where techniques like G-Blend [45] modulate gradients to combine unimodal and multimodal knowledge. OGM-GE [36] improves upon this by emphasizing modality-specific information, enhancing the model’s ability to leverage unique features from each input source. PMR [11] further refines this approach by modulating parameter groups based on the relevance of each modality, while QMF [54] introduces quadratic modulation for adaptive fusion. More recently, MLA [55] proposed alternating adaptation, which toggles between unimodal adaptation and gradient modulation to enable more efficient cross-modal learning. Another notable approach, MMLoRA [9], uses a two-phase strategy: unimodal fine-tuning followed by multimodal adaptation, which helps achieve optimal performance across a variety of tasks. However, while these methods allow for effective cross-modal learning, they often require significant computational resources and learning a large number of parameters to achieve optimal performance on downstream tasks.
@@
LoRA Rank $1$ $1$
LoRA Alpha $1$ $1$
@@
Image Resolution $224\times 224$ $224\times 224$
Patch Size $16\times 16$ $16\times 16$
@@
Num GPUs $2$ $1$
@@
Table 6: Performance comparison when modality gets missing during both training and inference. We train and evaluate our model with different modality ratio following [16, 20] and report the mean and standard deviation (SD) of three independent runs with different seeds.
Datasets Missing Rate ($\eta$) Train Inference ViLT MAP MSP Kim et al. CMPT (Ours)
Image Text Image Text [21] [22] [16] [20] (Mean $\pm$ SD)
@@
S2 Detailed Dataset Description
UPMC Food-101 dataset [46] is a popular and challenging multimodal classification dataset. It has two input modalities: image and text. The dataset has 90,704 samples divided into training and test sets having 67,988 and 22,716 samples, respectively. It contains 101 classes, which are the same as those in the ETHZ Food-101 dataset [3]. The samples are noisy, each category contains around 5% irrelevant images as they were collected in an uncontrolled environment without any human human intervention during the data collection process.
MM-IMDb dataset [2] is a widely used multimodal dataset for multi-label movie genre classification task. It consists of 25,959 samples, each containing both image and text as input modality. The dataset is divided into three subsets: 15,552 samples for training, 2,608 for validation, and 7,799 for testing. It has 23 classes and each sample can have one or more genre. This dataset serves as a standard benchmark for evaluating multimodal models in the multi-label classification task.
Kinetics-Sound (KS) dataset [1] is a subset of the Kinetics-400 dataset [17]. It is used for human action recognition using audio and video as input modalities. This dataset includes 31 different action classes that span a variety of everyday human activities, making it suitable for evaluating multimodal models. The dataset is split into two subsets: a training set containing 14,739 samples and a test set containing 2,594 samples.
Audio-Visual Event (AVE) Localization dataset [42] is a benchmark dataset used for multimodal event localization task. It contains 4,143 10-second videos. The dataset is divided into training, validation, and test sets, with 3,339 videos for training, 402 for validation, and 402 for testing. The dataset encompasses 28 different event categories providing a broad range of scenarios where both audio and visual signals are crucial for accurate event detection. All videos in the AVE dataset are collected from YouTube.
CREMA-D dataset [4] is used for multimodal emotion recognition. It has audio and video as input modalities. The dataset contains 7,442 short video clips performed by 91 actors. The clips cover six emotions: angry, happy, sad, neutral, disgust, and fear. Emotion labels were obtained from 2,443 crowd-sourced raters to ensure diverse evaluations. The dataset is divided into 6,698 samples for training and 744 samples for testing.
@@
Figure 6: Effectiveness of CMPT in handling missing modalities on UPMC Food-101 dataset. CMPT significantly improves performance when modalities are severely missing. Models are trained on 100% image + 100% text and tested with varying levels of missing modalities.
S3 Implementation Details
In Section 4.2 of the main paper, we briefly discussed the hyperparameter settings used in our experiments. Here, we provide a complete list of all the hyperparameters and their values in Table 5. The “Image-Text Datasets” column includes the hyperparameters for UPMC Food-101 and MM-IMDb datasets, while the “Audio-Video Datasets” column lists the hyperparameters for Kinetics-Sound, AVE, and CREMA-D datasets.
S4 Different Modality Ratios During Training and Inference
A robust multimodal learning method should maintain strong performance even when the missing modality ratio during inference differs from that during training. To evaluate the robustness of our approach, we follow the experimental setup of [16, 20] and train and test our model with varying modality ratios. Let $\eta$ denote the missing rate. We consider three scenarios: missing-text, missing-image, and missing-both. In the first two cases, $\eta\%$ of the samples contain only images or only text, while $(1-\eta)\%$ retain both modalities. For the missing-both case, $\frac{\eta}{2}\%$ of the samples contain only texts, $\frac{\eta}{2}\%$ contain only images, and $(1-\eta)\%$ include both modalities. We set $\eta=70$ throughout our experiments, following [16, 20].
As summarized in Table 6, our method consistently outperforms existing state-of-the-art approaches in most scenarios. While the method proposed by Kim et al. [20] performs well in certain cases, our approach demonstrates superior performance across both the MM-IMDb and UPMC Food-101 datasets. To ensure a rigorous evaluation, we report the mean and standard deviation (SD) over three independent runs with different seeds. The low SD values further indicate the stability and reliability of our method. These results substantiate our claim that CMPTs can effectively estimate missing modality features, enhancing robustness to varying missing modality conditions. Other methods reported in Table 1 do not report results for this specific experimental setup, and thus we are unable to compare with them.
S5 Effectiveness of CMPT
@@
We extend the study from Section 4.6.1 to evaluate the effectiveness of CMPT in handling missing modalities on the UPMC Food-101 dataset. We compare it against two baselines: (i) a standard model trained without modality dropout or CMPT (baseline) and (ii) a model trained with modality dropout, where modalities are randomly removed during training to improve robustness to missing data.
@@
As illustrated in Figure 6, the baseline model suffers a sharp performance decline as the proportion of missing modalities increases. While modality dropout improves robustness, it fails to fully recover the lost performance. Incorporating CMPT further enhances performance, achieving a 5.07% improvement over the baseline and 3.81% over modality dropout when the image modality is entirely missing as shown in Figure 6(a). A similar trend is observed when text is missing as shown in Figure 6(b)), where CMPT achieves 25.59% and 11.96% higher accuracy compared to the baseline and modality dropout, respectively. Consistent with our previous findings, CMPT becomes increasingly effective as the proportion of missing data increases, demonstrating its strong capability to mitigate performance degradation in the presence of severe modality loss.
@@
Figure 7: CMPT improves performance across most of the classes when text modality is missing on UPMC Food-101 dataset. However, it fails to accurately predict certain classes (e.g., bibimbap/sashimi), likely due to the image modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
@@
Figure 8: CMPT improves performance across most of the classes when image modality is missing on UPMC Food-101 dataset. However, it fails to accurately predict certain classes (e.g., lobster_roll_sandwitch/cheese_plate), likely due to the text modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
@@
S6 Does CMPT Benefit All Classes?
@@
As shown in Figure 7 and 8, CMPT improves or maintains performance across most of the classes in both text-missing and image-missing scenarios. Notably, it mitigates performance degradation in modality-sensitive classes. For instance, when text is missing, classes such as guacamole, pancakes, and foie_gras experience severe drops in performance, but CMPT effectively recovers much of the loss as shown in Figure 6(b). Similarly, in image-missing scenarios, CMPT provides substantial gains for sushi, cup_cakes, and pizza as shown in Figure 4(a). However, a slight performance drop is observed in certain classes when either text or image is absent, likely due to insufficient information in the available modality alone for those classes. Overall, these results reinforce our previous findings, demonstrating that CMPT enhances performance across most classes, leading to a strong overall performance improvement.
S7 Analysis of Modality Dependence

### v4

@@ 5 Limitations and Future Directions
While our approach demonstrates consistent performance improvements over existing methods in different missing modality scenarios, the remaining performance drop—especially in extreme cases—indicates room for further improvement. Additionally, we observe that certain classes are heavily dependent on a specific modality, which leads to significant performance drop in those classes when that specific modality is missing. A natural question arises as how can we design a better training strategy that encourages more uniform reliance on different modalities while maintaining overall performance in both complete and missing modality scenarios? Future work can also explore developing a general framework for explaining modality contribution and expected performance drop under different missing modality scenarios. Finally, this paper primarily focused on two-modality settings due to the availability of standard benchmarks and baselines that enable direct and fair comparisons. Extending the CMPT framework is a natural next step, and we are actively investigating this direction as part of our ongoing work.
@@ 6 Conclusion
We presented a simple yet effective framework for utilizing pre-trained unimodal encoders for robust multimodal learning. Our proposed framework enables training of multimodal systems that offer robust performance under different scenarios of missing modalities during training and inference. To compensate for missing modalities, we propose to learn cross-modal proxy tokens (CMPTs) to approximate the missing modality class tokens from the available modality using an alignment loss during training. Instead of training the unimodal encoders from scratch, we only learn rank-1 adapters along with task-specific classifier heads. Our extensive evaluations across diverse tasks and datasets demonstrate the robustness and adaptability of our approach in both complete and missing modality scenarios. Our analysis also reveals that CMPTs can generalize well to different missing scenarios and enhance performance across most of the classes to provide an overall performance boost. Overall, our approach provides a robust and efficient solution for multimodal learning that maintains high performance in different real-world scenarios.
Acknowledgment
This work is supported in part by NSF awards 2046293 and 2406199 and Amazon Gift awards.
Appendix A Appendix
Appendix A1 Related Work on Multimodal Learning
@@ Appendix A Appendix
In this section, we provide a general overview of multimodal learning, which involves integrating information from different input sources to improve performance on downstream tasks. We can broadly divide multimodal learning approaches into two main categories: designing specialized models and adapting existing models.
@@
Designing specialized models to effectively fuse multiple input sources is a common strategy in multimodal learning. For vision-language tasks, models such as ViLT (Kim et al., 2021), ViLBERT (Lu et al., 2019), and Visual-BERT (Li et al., 2019) process image patches and text tokens directly within a transformer architecture for tasks like image captioning and visual question answering. MMBT (Kiela et al., 2019) integrates both text and images within a BERT-based framework (Devlin et al., 2019) to model cross-modal dependencies. MBT (Nagrani et al., 2021b) leverages fusion bottlenecks for efficient modality-specific interaction, particularly in audio-video tasks. Specialized multimodal models have also been proposed for other applications, including action recognition (Woo et al., 2023; Zhou et al., 2022), segmentation (Zhang et al., 2023a; Wang et al., 2022b; Reza et al., 2024b), and sentiment analysis (Woo et al., 2023; Tsai et al., 2019). While these models excel in specific tasks, extending them to new modalities or domains often proves to be challenging and resource-intensive, as it requires retraining or significant modification of the model architecture.
@@
Adapting existing models offers an alternative approach to multimodal learning, focusing on the use of pre-trained unimodal models and adapting them for multimodal tasks. One such strategy is gradient modulation, where techniques like G-Blend (Wang et al., 2020a) modulate gradients to combine unimodal and multimodal knowledge. OGM-GE (Peng et al., 2022) improves upon this by emphasizing modality-specific information, enhancing the model’s ability to leverage unique features from each input source. PMR (Fan et al., 2023b) further refines this approach by modulating parameter groups based on the relevance of each modality, while QMF (Zhang et al., 2023b) introduces quadratic modulation for adaptive fusion. More recently, MLA (Zhang et al., 2024) proposed alternating adaptation, which toggles between unimodal adaptation and gradient modulation to enable more efficient cross-modal learning. Another notable approach, MMLoRA (Du et al., 2023), uses a two-phase strategy: unimodal fine-tuning followed by multimodal adaptation, which helps achieve optimal performance across a variety of tasks. However, while these methods allow for effective cross-modal learning, they often require significant computational resources and learning a large number of parameters to achieve optimal performance on downstream tasks.
@@
LoRA Rank 11 11
LoRA Alpha 11 11
@@
Image Resolution $\times 224 224×224$$\times 224$
Patch Size $\times 16 16×16$$\times 16$
@@
Num GPUs 22 11
Appendix A2 Detailed Dataset Description
UPMC Food-101 dataset (Wang et al., 2015) is a popular and challenging multimodal classification dataset. It has two input modalities: image and text. The dataset has 90,704 samples divided into training and test sets having 67,988 and 22,716 samples, respectively. It contains 101 classes, which are the same as those in the ETHZ Food-101 dataset (Bossard et al., 2014). The samples are noisy, each category contains around 5% irrelevant images as they were collected in an uncontrolled environment without any human human intervention during the data collection process.
MM-IMDb dataset (Arevalo et al., 2017) is a widely used multimodal dataset for multi-label movie genre classification task. It consists of 25,959 samples, each containing both image and text as input modality. The dataset is divided into three subsets: 15,552 samples for training, 2,608 for validation, and 7,799 for testing. It has 23 classes and each sample can have one or more genre. This dataset serves as a standard benchmark for evaluating multimodal models in the multi-label classification task.
Kinetics-Sound (KS) dataset (Arandjelovic & Zisserman, 2017) is a subset of the Kinetics-400 dataset (Kay et al., 2017). It is used for human action recognition using audio and video as input modalities. This dataset includes 31 different action classes that span a variety of everyday human activities, making it suitable for evaluating multimodal models. The dataset is split into two subsets: a training set containing 14,739 samples and a test set containing 2,594 samples.
Audio-Visual Event (AVE) Localization dataset (Tian et al., 2018) is a benchmark dataset used for multimodal event localization task. It contains 4,143 10-second videos. The dataset is divided into training, validation, and test sets, with 3,339 videos for training, 402 for validation, and 402 for testing. The dataset encompasses 28 different event categories providing a broad range of scenarios where both audio and visual signals are crucial for accurate event detection. All videos in the AVE dataset are collected from YouTube.
CREMA-D dataset (Cao et al., 2014) is used for multimodal emotion recognition. It has audio and video as input modalities. The dataset contains 7,442 short video clips performed by 91 actors. The clips cover six emotions: angry, happy, sad, neutral, disgust, and fear. Emotion labels were obtained from 2,443 crowd-sourced raters to ensure diverse evaluations. The dataset is divided into 6,698 samples for training and 744 samples for testing.
Table 7: Performance comparison with different LoRA ranks ($r$). The models are trained on complete multimodal data and tested on both complete and missing modality scenarios. The best Accuracy (Acc.) for each scenario is highlighted in bold. Params (M) indicates the number of learnable parameters in millions.
Datasets Available Modality Missing Modality r = 1 r = 2 r = 4
Params (M) Acc. (%) Params (M) Acc. (%) Params (M) Acc. (%)
UPMC Food-101 Image Text 0.271 75.66 0.376 75.62 0.671 70.38
Text Image 85.31 85.61 83.92
Image - Text - 94.47 94.64 93.30
AVE Audio Video 0.173 85.21 0.319 85.07 0.614 84.33
Video Audio 84.58 84.30 80.84
Audio - Video - 96.77 96.77 96.27
CREMA-D Audio Video 0.157 67.20 0.303 65.32 0.598 65.46
Video Audio 76.21 75.54 76.61
Audio - Video - 88.84 87.63 87.77
KS Audio Video 0.176 68.27 0.322 69.93 0.617 69.54
Video Audio 85.77 85.20 86.43
Audio - Video - 91.21 91.06 90.59
Appendix A3 Implementation Details
In Section 4.2 of the main paper, we briefly discussed the hyperparameter settings used in our experiments. Here, we provide a complete list of all the hyperparameters and their values in Table 6. The “Image-Text Datasets" column includes the hyperparameters for UPMC Food-101 and MM-IMDb datasets, while the “Audio-Video Datasets" column lists the hyperparameters for Kinetics-Sound, AVE, and CREMA-D datasets.
Appendix A4 Determining the Optimal LoRA Rank
To determine the optimal LoRA rank, we conducted an ablation study evaluating ranks $r\in\{1,2,4\}$. For this analysis, we trained our models on modality-complete data and evaluated their performance in both complete and missing modality scenarios, with the results presented in Table 7.
Our analysis indicates that increasing the rank, and thereby the number of learnable parameters, does not guarantee a consistent improvement in performance. For instance, on the AVE dataset, a rank of $r=1$ consistently achieves the highest accuracy. On the UPMC Food-101 dataset, performance between $r=1$ and $r=2$ is comparable. While $r=2$ offers a marginal improvement in two scenarios, it requires a 38.7% increase in learnable parameters. In contrast, using $r=4$ consistently resulted in degraded performance in both of these datasets compared to $r=1$. A similar trend is observed for the CREMA-D dataset, where $r=1$ performs best in two of the three scenarios. Conversely, the KS dataset presented mixed results, with different ranks yielding optimal performance in different scenarios.
Overall, a rank of $r=1$ demonstrates competitive, and often superior performance across a majority of the tested scenarios. The performance degradation at higher ranks, such as $r=4$, suggest that a larger number of learnable parameters may not be necessary for these tasks and could introduce a risk of overfitting. This aligns with the common characteristic of LoRA, where a low rank is often sufficient to capture essential task-specific information. Based on these observations, we use LoRA with $r=1$ for all our experiments as it provides an optimal balance, delivering robust performance with the greatest parameter efficiency.
@@
Table 8: Performance comparison when modality gets missing during both training and inference. We train and evaluate our model with different modality ratio following existing baselines (Jang et al., 2024; Kim & Kim, 2024) and report the mean and standard deviation (SD) of three independent runs with different seeds.
Datasets Missing Train Inference ViLT MAP MSP Kim & Kim CMPT (Ours)
Rate ($\eta$) Image Text Image Text (2021) (2023) (2024) (2024) (Mean $\pm$ SD)
@@
Appendix A5 Different Modality Ratios During Training and Inference
A robust multimodal learning method should maintain strong performance even when the missing modality ratio during inference differs from that during training. To evaluate the robustness of our approach, we follow the experimental setup of Jang et al. (2024); Kim & Kim (2024) and train and test our model with varying modality ratios. Let $\eta$ denote the missing rate. We consider three scenarios: missing-text, missing-image, and missing-both. In the first two cases, $\eta\%$ of the samples contain only images or only text, while $(1-\eta)\%$ retain both modalities. For the missing-both case, $\frac{\eta}{2}\%$ of the samples contain only texts, $\frac{\eta}{2}\%$ contain only images, and $(1-\eta)\%$ include both modalities. We set $\eta=70$ throughout our experiments, following existing baselines (Jang et al., 2024; Kim & Kim, 2024).
As summarized in Table 8, our method consistently outperforms existing state-of-the-art approaches in most scenarios. While the method proposed by Kim & Kim (2024) performs well in certain cases, our approach demonstrates superior performance across both the MM-IMDb and UPMC Food-101 datasets. To ensure a rigorous evaluation, we report the mean and standard deviation (SD) over three independent runs with different seeds. The low SD values further indicate the stability and reliability of our method. These results substantiate our claim that CMPTs can effectively estimate missing modality features, enhancing robustness to varying missing modality conditions. Other methods reported in Table 1 do not report results for this specific experimental setup, and thus we are unable to compare with them.
@@
Figure 7: Effectiveness of CMPTs in handling missing modalities on UPMC Food-101 dataset. CMPTs significantly improves performance when modalities are severely missing. Models are trained on 100% image + 100% text and tested with varying levels of missing modalities.
Appendix A6 Effectiveness of CMPTs
@@
We extend the study from Section 4.6.1 to evaluate the effectiveness of CMPTs in handling missing modalities on the UPMC Food-101 dataset. We compare it against two baselines: (i) a standard model trained without modality dropout or CMPTs (baseline) and (ii) a model trained with modality dropout, where modalities are randomly removed during training to improve robustness to missing data. For a fair comparison, all models were fine-tuned on the downstream task using LoRA with a rank of 1.
@@
As illustrated in Figure 7, the baseline model experiences a significant performance drop as the proportion of missing modalities increases. While modality dropout improves robustness, it fails to fully recover the lost performance. Incorporating CMPTs further enhances performance, achieving a 5.07% improvement over the baseline and 3.81% over modality dropout when the image modality is entirely missing as shown in Figure 7(a). A similar trend is observed when text is missing as shown in Figure 7(b), where CMPTs achieves 25.59% and 11.96% higher accuracy compared to the baseline and modality dropout, respectively. Consistent with our previous findings, CMPTs becomes increasingly effective as the proportion of missing data increases, demonstrating its strong capability to substantially mitigate performance degradation in the presence of severe modality loss.
@@
Figure 8: CMPTs improve performance across most of the classes when text modality is missing on UPMC Food-101 dataset. However, it fails to accurately predict certain classes (e.g., bibimbap/sashimi), likely due to the image modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
@@
Figure 9: CMPTs improve performance across most of the classes when image modality is missing on UPMC Food-101 dataset. However, it fails to accurately predict certain classes (e.g., lobster_roll_sandwitch/cheese_plate), likely due to the text modality’s insufficient information for those classes. Classes are sorted based on the amount of performance improvement.
Figure 10: t-SNE plots of the fused feature tokens $\mathcal{T}$ for ten classes where the performance gains from CMPTs are limited. In these cases, the fused feature embeddings with CMPTs (blue) remain diffuse and overlapping, and exhibit weak alignment with the complete multimodal features (green). Although CMPTs show improvement over the baseline without CMPTs (red), the alignment remains insufficient to form consistently discriminative clusters, indicating limited or no improvement in missing modality robustness in these classes.
Figure 11: Attention map visualizations from the last layer of the vision encoder for the CLS token and CMPTs in the text-missing scenario on the UPMC Food-101 dataset. Left: Cases where the fused features with and without CMPTs provide correct predictions. Here, CLS tokens already attend to semantically-relevant regions and CMPTs provide further refinement, but do not change the prediction. Right: Cases where CMPTs attend to semantically-relevant regions but lead to incorrect predictions. The last example illustrates a scenario where both the fused features with and without CMPTs result in incorrect predictions. These examples highlight limitations of CMPTs when the available modality lacks sufficient discriminative information or when learned cross-modal associations are imprecise.
@@
Appendix A7 Do CMPTs Benefit All Classes?
@@
As shown in Figure 8 and 9, CMPTs improve or maintain performance across most of the classes in both text-missing and image-missing scenarios. Notably, it mitigates performance degradation in modality-sensitive classes. For instance, when text is missing, classes such as guacamole, pancakes, and foie_gras experience severe drops in performance, but CMPTs effectively recovers much of the loss as shown in Figure 7(b). Similarly, in image-missing scenarios, CMPTs provide substantial gains for sushi, cup_cakes, and pizza as shown in Figure 3(a). However, a slight performance drop is observed in certain classes when either text or image is absent. For instance, CMPT fails to recover performance for classes such as bibimbap and sashimi in the absence of text, and for classes like lobster_roll_sandwich and cheese_plate when the image modality is missing. This is likely due to insufficient information in the available modality alone for those classes. Overall, these results reinforce our previous findings, demonstrating that CMPTs enhance performance across most classes, leading to a strong overall performance improvement.
Appendix A8 Limitations and Failure Case Analysis
While CMPTs demonstrate notable improvements in robustness, it is essential to understand their limitations. Figure 10 presents t-SNE plots of the fused feature tokens $\mathcal{T}$ for ten classes on the UPMC Food-101 dataset where the performance gains from CMPTs are limited. In these cases, the fused feature embeddings with CMPTs (blue) remain diffuse and overlapping, and exhibit weak alignment with the complete multimodal features (green). This is particularly evident for visually ambiguous classes or classes with high intra-class variance in the feature embedding space. Although CMPTs show improvement over the baseline without CMPTs (red), the alignment remains insufficient to form consistently discriminative clusters, indicating limited or no improvement in missing modality robustness.
We also analyze the attention maps from the last layer of the vision encoder for the CLS token and CMPTs in the text-missing scenario, as shown in Figure 11. The left half of the figure includes examples when the fused features with and without CMPTs provide correct predictions (approximately 60% samples show this pattern). For these cases, the CLS token already attends to semantically-relevant regions and make the correct prediction. CMPTs further refine the attention but do not change the model’s prediction, as it is already correct. Thus, the inclusion of CMPTs does not necessarily improve performance. The right half of the figure shows examples when the fused features with CMPTs lead to incorrect predictions (approximately 3% samples show this pattern). We observe that CMPTs attend to semantically-relevant regions of the image but still lead to incorrect predictions. For example, in the French Onion Soup sample, the CMPT focuses heavily on the melted cheese surface—a visually salient but ambiguous feature—leading to a misclassification as Shrimp and Grits. Similarly, for the Lobster Roll Sandwich, the CMPT fails to recover from the original misclassification and predicts Foie Gras based on misleading visual cues. These examples highlight scenarios where either the available modality is insufficient or the learned cross-modal associations are misleading, resulting in CMPTs contributing to incorrect predictions rather than correcting them.
Appendix A9 Analysis of Modality Dependence
@@
Table 9: Performance comparison with three different parameter efficient adaptation methods. We used rank=1 for all the methods. The best Accuracy (Acc.) for each scenario is highlighted in bold. Params (M) indicates the number of learnable parameters in millions.
Datasets Available Modality Missing Modality LoRA DoRA VeRA
Params (M) Acc. (%) Params (M) Acc. (%) Params (M) Acc. (%)
UPMC Food-101 Image Text 0.271 75.66 0.302 69.03 0.155 70.82
Text Image 85.31 83.52 82.85
Image - Text - 94.47 92.59 92.34
AVE Audio Video 0.173 85.21 0.246 86.57 0.098 89.05
Video Audio 84.58 91.79 90.05
Audio - Video - 96.77 96.27 96.27
CREMA-D Audio Video 0.157 67.20 0.229 67.07 0.082 66.13
Video Audio 76.21 73.12 71.64
Audio - Video - 88.84 88.04 86.02
KS Audio Video 0.176 68.27 0.248 68.77 0.101 69.93
Video Audio 85.77 85.97 83.69
Audio - Video - 91.21 91.32 90.71
Appendix A10 Generalizability to Other Parameter-Efficient Adaptation Methods
To demonstrate that CMPTs are not restricted to a particular parameter-efficient adaptation method, we extend our analysis beyond LoRA (Hu et al., 2022), which is used in our primary experiments. Specifically, we evaluate CMPTs with two recent alternatives: DoRA (Liu et al., 2024) and VeRA (Kopiczko et al., 2024). For a fair comparison, we fix the rank to $1$ across all methods. We train the models on modality-complete data and evaluate them on both modality-complete and modality-incomplete scenarios. The results are summarized in Table 9.
We find that the number of learnable parameters differs across methods, with VeRA being the most parameter-efficient and DoRA the least. In terms of performance, no single method consistently outperforms the others across all datasets. LoRA achieves the best overall results on the UPMC Food-101 and CREMA-D datasets, while DoRA shows superior performance on the Kinetics-Sound dataset and remains highly competitive on AVE. VeRA, while using substantially fewer parameters, delivers strong performance but is generally slightly below LoRA and DoRA, suggesting a trade-off between parameter efficiency and maximal performance.
Most importantly, our framework maintains robust performance in missing-modality scenarios regardless of the underlying adaptation method. The relative stability observed between complete and missing-modality settings is consistent across all three adaptation techniques. These findings confirm that CMPTs function as a modular approach whose effectiveness is not tied to any specific adaptation strategy, underscoring their scalability and versatility across different parameter-efficient methods.
Table 10: Performance comparison of our method with a ViT versus a SigLIP model as the vision backbone. The models are trained on complete multimodal data and tested on both complete and missing modality scenarios. We report % accuracy as the performance metric, and the highest score for each scenario is highlighted in bold.
Datasets Available Modality Missing Modality ViT SigLIP
(Dosovitskiy et al. (2021)) (Zhai et al. (2023))
UPMC Food-101 Image Text 75.66 77.50
Text Image 85.31 85.58
Image - Text - 94.47 94.58
AVE Audio Video 85.21 86.32
Video Audio 84.58 30.60
Audio - Video - 96.77 94.78
CREMA-D Audio Video 67.20 65.73
Video Audio 76.21 35.35
Audio - Video - 88.84 84.68
KS Audio Video 68.27 69.66
Video Audio 85.77 30.76
Audio - Video - 91.21 89.44
Table 11: Performance comparison of our method using Base versus Large ViT and BERT models. The models are trained on complete multimodal data and tested on both complete and missing modality scenarios. We report % accuracy for UPMC Food-101 and F1-Macro score for MM-IMDb dataset. Params (M) indicates the number of learnable parameters in millions. The best performance for each scenario is highlighted in bold.
Datasets Available Missing ViT-Base and BERT-Base ViT-Large and BERT-Large
Modality Modality Params (M) Performance Params (M) Performance
MM-IMDb (F1-Macro) Image Text 0.211 46.97 0.540 53.42
Text Image 56.32 59.92
Image - Text - 63.58 64.00
UPMC Food-101 (Accuracy) Image Text 0.271 75.66 0.501 77.17
Text Image 85.31 83.91
Image - Text - 94.47 93.74
Appendix A11 Applicability to SigLIP-Based Models
To assess the architectural flexibility of our method, we replaced the ViT encoder (Dosovitskiy et al., 2021) with a pretrained SigLIP model (Zhai et al., 2023), which employs Global Attention Pooling (GAP) instead of a CLS token. For audio–video datasets, we retained the AST model as the audio encoder, and for image–text datasets, we used BERT as the text encoder. In this setup, CMPT tokens must perform a cross-architectural reconstruction: inferring missing audio or text features from SigLIP’s vision embedding, and vice versa.
We train the models on modality-complete data and evaluate them on both modality-complete and modality-incomplete scenarios. The results in Table 10 show that our method is not restricted to a single encoder type. On UPMC Food-101 dataset, the SigLIP-based model consistently outperforms the ViT-based model, demonstrating that CMPTs can be effectively applied to GAP-based vision backbones for vision–language tasks. In contrast, for audio–video datasets, SigLIP exhibits sharp performance degradation when audio is missing. Here, CMPTs struggle to recover audio features from global video embeddings, likely because GAP discards the fine-grained temporal information that is critical for inferring a corresponding audio representation. The reverse process, however, is more successful. When the video modality is missing, AST’s audio features provide a strong basis for reconstructing SigLIP’s global video representation. This suggests that AST’s feature space is more compatible with SigLIP than vice versa.
Overall, these findings indicate that CMPTs remain applicable to GAP-based models like SigLIP, but their effectiveness depends on the modality. While GAP facilitates efficient vision embeddings, it limits the representational richness required for audio reconstruction, constraining CMPTs in certain cases. Nonetheless, the results confirm that CMPTs are not inherently incompatible with SigLIP, though their utility is shaped by architectural and modality-specific factors.
Appendix A12 Scalability with Larger Vision and Text Encoders
To assess the scalability of our framework with larger encoders, we replaced the ViT-Base and BERT-Base encoders used in our primary experiments with ViT-Large and BERT-Large. We train the models on modality-complete data and evaluate them on both modality-complete and modality-incomplete scenarios on the UPMC Food-101 and MM-IMDb datasets.
The results, summarized in Table 11, show that our framework can leverage the increased capacity of larger backbones. On MM-IMDb, scaling yields consistent improvements in F1-Macro across all scenarios. The gains are particularly notable in missing-modality cases, with a +6.45% increase when text modality is missing and +3.60% when image modality is missing. These results suggest that stronger backbone representations improve CMPT’s ability to reconstruct missing features and maintain robust predictions.
On UPMC Food-101, the larger backbone improves image-only accuracy by +1.51% but shows a decline of -1.41% for text-only performance. With both modalities available, performance remains comparable (94.47 vs. 93.74). This indicates a potential trade-off, where the added capacity may overfit to the text modality, reducing generalization when the image is absent.
In summary, CMPT scales effectively with larger backbones, particularly on complex datasets such as MM-IMDb. At the same time, the results highlight that the benefits of scaling depend on dataset and modality characteristics, underscoring the need for careful backbone selection to balance capacity and generalization.